OpenStudy (anonymous):

What are the x-intercepts of y = 2x^2 - 3x - 20?

OpenStudy (campbell_st):

factorise the quadratic (x-4) (2x-5) = 0 then solve: x - 4 = 0 2x + 5 = 0 to find thex intercepts

OpenStudy (anonymous):

A) x = -5 and 2 B) x = 5 and -2 C) x = -2.5 and 4 D) x = 2.5 and -4

OpenStudy (anonymous):

c?

OpenStudy (campbell_st):

looks like it

OpenStudy (anonymous):

thanks